Steam generated in a Rankine steam power cycle at a pressure of 8 MPa and temperature 600oC is fed to a turbine. Exhaust from the turbine enters a condenser at 100 kPa, where it is condensed to saturated liquid, which is then pumped to the boiler. (a) (20 pts.) What is the thermal efficiency η of an ideal Rankine cycle operating at these conditions? (b) (5 pts.) If the net power production of the cycle is 80,000 kW, what is the steam mass flow rate [in kg/s]? (c) (15 pts.) What is the thermal efficiency of the power cycle if the turbine efficiency and pump efficiency are 75%
Answer #1

Given Data - Boiler pressure P1 = Py = 8 MPa Condenser pressure P2 = P3 = 100 kPa Turbine inlet Ti = Gooc solution T T-S Diag• h₂= hf+42x hfg @ look la @ look pa - 417.5+0.9434 2257.9 h2 = 2548.28 83 kJ/kg hz=hf 417.5 kJ kJ/kg @look la = O pump workWnet = 108 2.97 kJ/kg a thermal efficiency na Whet = Qs 1082.97.= 0.3369 3213.76 33.6 Aus y = 33.69% Aus b Net power = 80,000Up puru efficiency hy-43 hy-h3 3 0.75= 425.74 - 417.5 417.5 hy. hy= 428.406 kJ/kg Actual turbine work => Wq ha ha ha ha = 81